### v1

# Making it to First: The Random Access Problem in DNA Storage

Abstract
We study the Random Access Problem in DNA storage, which addresses the challenge of retrieving a specific information strand from a DNA-based storage system. Given that $k$ information strands, representing the data, are encoded into $n$ strands using a code. The goal under this paradigm is to identify and analyze codes that minimize the expected number of reads required to retrieve any of the $k$ information strand, while in each read one of the $n$ encoded strands is read uniformly at random. We fully solve the case when $k=2$, showing that the best possible code attains a random access expectation of $0.914\cdot 2$. Moreover, we generalize a construction from [17], specific to $k=3$, for any value of $k$. Our construction uses $B_{k-1}$ sequences over $\mathbb{Z}_{q-1}$, that always exist over large finite fields. For $k=4$, we show that this generalized construction outperforms all previous constructions in terms of reducing the random access expectation.
00footnotetext: A. B. and E. Y. are supported by the European Union (DiDAX, 101115134). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. O. E. is supported by the Israel Science Foundation (Grant No. 1789/23). R. G. is supported by NSF Grant CCF2212437. A. G. is supported by the Villum Fonden Grant VIL”52303”. I. T. is supported by the European Research Council (Grant No. 852953). Email addresses: {avital.bor,yaakobi}@cs.technion.ac.il, ohadeli@bgu.ac.il, rgabrys@ucsd.edu, anigr@dtu.dk, tamo@tauex.tau.ac.il
1 Introduction
The exponential growth in data generation has created an unprecedented demand for storage technologies, which current solutions are unable to meet. The gap between data storage demand and the capacity of existing technologies continues to widen at an alarming rate each year [25]. Addressing this critical challenge has become a global priority, driving the search for innovative and sustainable alternatives. One particularly promising approach is the use of synthetic DNA as a medium for data storage [10, 20].
A typical DNA data storage system consists of three primary components: DNA synthesis, storage containers, and DNA sequencing. In the first step, synthetic DNA strands, known as oligos, are generated to encode the user’s information. These strands are then stored in an unordered manner within a storage container. In the final step, DNA sequencing reads the stored strands and converts them into digital sequences, referred to as reads, which are decoded back into the original user information. However, due to limitations in current technologies, the process produces multiple noisy copies of each designed strand, and these copies are retrieved in a completely unordered fashion.
While several studies have demonstrated the significant potential of DNA as a data storage medium [2, 6, 7, 22, 27, 28, 3], its adoption as a practical alternative to current storage technologies is still limited by challenges related to cost and efficiency. A key factor contributing to these challenges is the coverage depth of DNA storage, defined as the ratio between the number of sequenced reads and the number of designed strands[18]. Reducing the coverage depth is critical for improving the latency of existing DNA storage systems and significantly lowering their costs. One very natural question concerning the coverage depth has been studied recently in [4] and deals with the problem of reducing the expected number of samples needed to retrieve all the stored information. This problem is closely related to the Coupon Collector’s Problem [13, 14, 12, 21], for more details, see the last paragraph in Section 2.
In this work, we study the Random Access Problem, first introduced in [4], which addresses the challenge of retrieving a single specific information strand from a DNA-based storage system. Given that $k$ information strands, representing the data, are encoded into $n$ strands using an error-correcting code, the goal is to identify and analyze codes that minimize the expected number of reads required to retrieve any of the $k$ information strands, while in each read one of the $n$ encoded strands is read uniformly at random. Initial steps towards solving this problem were made in [4] and [16], where constructions, analysis of well-known codes, and bounds on the random access expectation for arbitrary codes were presented. In a later work [17], this setting was explored from a geometric perspective, leading to a construction that outperformed all previously known codes for $k=3$. Despite the progress made in these works, many essential questions remain unsolved.
In this paper, we address some of the open questions in the random access setting. In particular, we provide a complete solution for the $k=2$ case, showing what the smallest possible random access expectation is in this case. Moreover, we generalize the construction from [17] for arbitrary values of $k$ and show the resulting construction outperforms all current known codes for $k=4$. The rest of the paper is organized as follows. Section 2 introduces the relevant definitions and formalizes the problem statement. Section 3 provides a solution to the random access problem for $k=2$. In Section 4, we provide a construction for general $k$ and then analyze it for $k=3$ and $k=4$. Finally, Section 5 concludes the paper and outlines several directions for future research.
2 Preliminaries
Throughout this paper, we adopt the following notations. Let $k$ and $n$ be positive integers with $k\leq n$, and let $q$ be a prime power. Denote by $[k]$ the set $\{1,2,\dots,k\}$ and by $H_{n}$ the $n$-th harmonic number, i.e., $H_{n}:=1+1/2+\dots+1/n$. Let $\mathbb{F}_{q}$ denote the finite field with $q$ elements, and let ${\mathbb{F}_{q}^{k}}$ denote the $k$-dimensional vector space over ${\mathbb{F}}_{q}$. Let ${{\boldsymbol{e}}}_{i}\in{\mathbb{F}}_{q}^{k}$ denote the $i$-th standard vector, and for a set of vectors ${\bf g}_{1},{\bf g}_{2},\dots,{\bf g}_{i}\in{\mathbb{F}}_{q}^{k}$, let $\langle{\bf g}_{1},{\bf g}_{2},\dots,{\bf g}_{i}\rangle$ denote their span.
We study the expected sample size required for uniform random access in DNA storage systems. In such systems, data is stored as a length-$k$ vector of sequences, referred to as strands, each of length $\ell$ over the alphabet $\Sigma=\{A,C,G,T\}$. This corresponds to representing data as elements in $(\Sigma^{\ell})^{k}$. We can embed $\Sigma^{\ell}$ into a finite field $\mathbb{F}_{q}$, which requires $4^{\ell}$ to divide $q$. However, in this work, we study this problem in a more general setting, considering any prime power $q$ for the size of our underlying finite field.
The encoding process utilizes a $k$-dimensional linear block code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$, mapping an information vector ${\boldsymbol{x}}=(x_{1},\dots,x_{k})\in\mathbb{F}_{q}^{k}$ to an encoded vector ${\boldsymbol{y}}=(y_{1},\dots,y_{n})\in\mathbb{F}_{q}^{n}$. To retrieve the stored information, the encoded strands are first synthesized and then sequenced using DNA sequencing technology. This process produces multiple erroneous copies of the strands, referred to as reads. In line with prior works [4, 16, 17], we assume that no errors are introduced during synthesis or sequencing. The output of the sequencing process is thus a multiset of unordered reads. Given the high cost and relatively low throughput of current DNA sequencing technologies compared to other archival storage systems, reducing the coverage depth—the ratio of sequenced reads to the number of encoded strands—is crucial for improving efficiency.
In this paper, we focus on the random access setting, where the goal is to retrieve a single information strand $x_{i}$ for $i\in[k]$. Previous works (e.g., [4, 16, 17]), have shown that appropriate coding schemes can reduce the expected sample size for recovering an information strand to below $k$. We illustrate this in Example 1, as presented in [16]. Note that we implicitly assume that we know the index of each of the strands that we are sampling. If we did not know the locations of the recovered symbols in the underlying code, we do not necessarily know how to recover information strands from sampled strands.
Example 1.
Assume we want to store an information vector of size two, $(x_{1},x_{2})\in{\mathbb{F}}_{q}^{2}$. Without coding, the expected number of samples required to recover a specific information strand is 2, assuming that the samples are read uniformly at random. Now, suppose we encode the data using the following generator matrix $G$
$G=\begin{pmatrix}1&0&1&0&1\\ 0&1&0&1&1\\ \end{pmatrix}.$ (1)
The encoded data is stored as $(x_{1},x_{2})G=(x_{1},x_{2},x_{1},x_{2},x_{1}+x_{2})\in{\mathbb{F}}_{q}^{5}$. Using the known expectation of a geometric distribution and conditioning on the first sample, it can be shown that the expected number of samples required to recover a specific information strand for this case is approximately $1.917<2$. Here, “recovering” means reconstructing the original information strand as a linear combination of the sampled symbols. For instance, if the last two encoded symbols are sampled, the strand $x_{1}$ can be recovered as $x_{1}=x_{2}+(x_{1}+x_{2})$.
Example 1 demonstrates that once the $k$ information strands are encoded using a generator matrix $G\in{\mathbb{F}}_{q}^{k\times n}$, each encoded strand corresponds to a column of $G$. Moreover, recovering the $i$-th information strand is equivalent to sampling a collection of columns for which the $i$-th standard basis vector, ${{\boldsymbol{e}}}_{i}$, lies in their ${\mathbb{F}}_{q}$-span.
Remark 1.
Because of the discussion above, we only care about the ${\mathbb{F}}_{q}$-span of the sampled columns of the generator matrix $G$. This means in particular, that the order in which the columns show up in the matrix does not matter, and replacing any column by a collinear column will result in the same expected number of samples to recover any information strand.
To formalize the problem, we define the main problem addressed in this paper, building on the framework presented in [16].
Problem 1.
Let $G\in{\mathbb{F}}_{q}^{k\times n}$ be a rank-$k$ matrix. Suppose that the columns of $G$ are drawn uniformly at random, meaning that each column has a probability $1/n$ of being drawn and columns can be drawn multiple times. For $i\in[k]$, let $\tau_{i}(G)$ denote the random variable that counts the minimum number of columns of $G$ that are drawn until the standard basis vector ${{\boldsymbol{e}}}_{i}\in{\mathbb{F}}_{q}^{k}$ is in their ${\mathbb{F}}_{q}$-span. Our goals are:
  1. 1.
    Compute the expectation ${\mathbb{E}}[\tau_{i}(G)]$ and the maximum expectation:
    $T_{\max}(G)=\max_{i\in[k]}{\mathbb{E}}[\tau_{i}(G)].$
  2. 2.
    Determine the smallest possible maximum random access expectation among all rank-$k$ matrices:
    $T_{q}(n,k)=\min_{G\in{\mathbb{F}}_{q}^{k\times n}}T_{\max}(G),$
    the asymptotic behavior as $n$ approaches infinity:
    $T_{q}(k)=\liminf_{n\longrightarrow\infty}T_{q}(n,k),$
    and as $q$ approaches infinity:
    $T(k)=\liminf_{q\longrightarrow\infty}T_{q}(k).$
Note that the values of $T_{q}(n,k)$ and $T_{q}(k)$ are defined only when $q$ is a prime power. If $Q$ is the set of all prime powers, we omit writing $q\in Q$ when writing $q\to+\infty$.
The study of the values of $T_{\max}(G),\text{\leavevmode\nobreak\ }T_{q}(n,k),\text{\leavevmode\nobreak\ }T_{q}(k),$ and $T(k)$ was initiated in [4]. It was shown that for several families of codes, such as the identity code, the simple parity code, and MDS codes, it holds that ${\mathbb{E}}[\tau_{i}(G)]=k$ for every $i\in[k]$, when $G$ is a systematic generator matrix of these codes. In particular, the result for identity codes established that $T_{q}(k,k)=k$. However, determining $T_{q}(n,k)$ for general parameters remains an open and intriguing question. Initial progress towards addressing this problem was made in [4], where several constructions were proposed that achieve a maximum expectation strictly less than $k$. In particular, it was shown that $T(k=2)\leq 0.914\cdot 2,\text{\leavevmode\nobreak\ }T(k=3)\leq 0.89\cdot 3$ and for arbitrary $k$ which is a multiple of $4$, it holds that for large enough $q$: $T_{q}(n=2k,k)\leq 0.95k$. In addition, two lower bounds on $T_{q}(n,k)$ were derived in [4]: for any $n,k$ and $q$, we have that $T_{q}(n,k)\geq n-\frac{n(n-k)}{n}(H_{n}-H_{n-k})$ and $T_{q}(n,k)\geq\frac{k+1}{2}$.
A significant challenge in the random access problem lies in the difficulty of directly computing $T_{\max}(G)$. To address this, the authors of [16] provided a general formula for the expected number of reads required to recover the $i$-th information strand.
Lemma 1 (see [16], Lemma 1).
For $G\in{\mathbb{F}}_{q}^{k\times n}$, let
$\alpha_{i}^{s}(G)=|\{S\subseteq[n]:|S|=s,\text{\leavevmode\nobreak\ }{{\boldsymbol{e}}}_{i}\in\langle{{\boldsymbol{g}}}_{j}:\text{\leavevmode\nobreak\ }j\in S\rangle\},$
where ${{\boldsymbol{g}}}_{j}$ represents the $j$-th column of $G$. Then, for every $i\in[k]$, the expected value of $\tau_{i}(G)$ is given by:
${\mathbb{E}}[\tau_{i}(G)]=nH_{n}-\sum_{s=1}^{n-1}\frac{\alpha_{i}^{s}(G)}{{n-1\choose s}}.$ (2)
Using this result, it was derived that various families of codes, e.g. Hamming, simplex, systematic MDS, all achieve a random access expectation of $k$. Then in [17], by looking at the random access problem from a geometric point of view, the authors proposed a construction for the case of $k=3$. Using the above formula, they demonstrated that their construction achieves an expectation upper bounded by $0.88\overline{22}\cdot 3$, improving the result of [4]. In addition, [17] resolved a conjecture proposed in [4] regarding a class of codes with rate $1/2$, which achieves a random access expectation strictly smaller than $0.9456k$ as $k$ tends to infinity.
Despite these valuable contributions, the fundamental limits of the random access coverage depth problem are still not well understood. Even for the case of $k=2$, the values of $T_{q}(k)$ and $T(k)$ had not been determined previously. In this work, we seek to deepen our understanding of the random access problem and provide new insights into the values of $T_{\max}(G),\text{\leavevmode\nobreak\ }T_{q}(n,k),\text{\leavevmode\nobreak\ }T_{q}(k)$ and $T(k)$.
We would like to note that a closely related problem is the non-random access setting, where the goal is to retrieve all the user’s information. By drawing a connection to the Coupon Collector’s Problem [13, 14, 12, 21], the authors in [4] established that if the $k$ information strands are encoded by an MDS code, then the expected number of reads to decode all $k$ information strands is $H_{n}-H_{n-k}$. Moreover, the authors showed that one cannot do better than this, and so MDS codes are the most suitable codes for minimizing the expected number of reads when the user wants to retrieve all their data. Later on, the non-random access version of the DNA coverage depth problem was expanded in [9] to address composite DNA letters [2], and further explored in [23, 26] for the setup of the combinatorial composite of DNA shortmers [24]. A similar concept was examined in [8], where the authors analyzed the trade-offs between reading costs, tied directly to coverage depth, and writing costs. Another extension to the random access setup was studied in [1]. However, their study focused on decoding groups of DNA strands collectively representing a single file, rather than decoding individual strands.
3 Optimal Random Access Expectation for Two Information Strands
In this section, we determine the exact values of $T_{q}(2)$ and $T(2)$. We start by identifying the structure of a $q$-ary two-row matrix $G$ that minimizes $T_{\max}(G)$.
Notation 1.
In this section, let $x_{1}$ and $x_{2}$ denote the number of columns in $G$ of the form $(1,0)^{T}$ and $(0,1)^{T}$ respectively (recall that these vectors correspond to the information strands). Observe that over ${\mathbb{F}}_{q}$ there are exactly $q-1$ vectors distinct from the information strands such that no two distinct vectors among them are collinear; the only columns we need to be concerned about are those that are not collinear because of Remark 1. Denoting by $\beta\in{\mathbb{F}}_{q}\setminus\{0\}$ a primitive element, these vectors can be expressed as $(1,\beta^{i})^{T}$ for $20\leq i\leq q-$ 2. Let $a_{i}$ denote the number of columns in $G$ of the form $(1,\beta^{i})^{T}$, then the total number of columns in $G$ is given by $x\triangleq x_{1}+x_{2}+\sum_{i=0}^{q-2}a_{i}$.
In the next claim we calculate the expectation of $\tau_{1}(G)$ and $\tau_{2}(G)$.
Claim 1.
It holds that:
${\mathbb{E}}[\tau_{1}(G)]=1+\frac{x_{2}}{x-x_{2}}+\sum_{i=0}^{q-2}\frac{a_{i}}{x-a_{i}},$ (3)
${\mathbb{E}}[\tau_{2}(G)]=1+\frac{x_{1}}{x-x_{1}}+\sum_{i=0}^{q-2}\frac{a_{i}}{x-a_{i}}.$ (4)
Proof.
We condition on the first draw. If the first draw yields a column of the form $(0,1)^{T}$, then, upon subsequently drawing a column of a different form, the first column can be recovered. Consequently, given that the first sample was $(0,1)^{T}$, $\tau_{1}(G)$ has a geometric distribution with success probability ${(x-x_{2})}/{x}$. Analogous reasoning applies if the first draw corresponds to a column of the form $(1,\beta^{i})^{T}$. Therefore, the expected waiting time ${\mathbb{E}}[\tau_{1}(G)]$ is given by
${\mathbb{E}}[\tau_{1}(G)]$ $=\frac{x_{1}}{x}+\frac{x_{2}}{x}(1+\frac{x}{x-x_{2}})+\sum_{i=0}^{q-2}\frac{a_{i}}{x}(1+\frac{x}{x-a_{i}})$
$=1+\frac{x_{2}}{x-x_{2}}+\sum_{i=0}^{q-2}\frac{a_{i}}{x-a_{i}}.$
$E[\tau_{2}(G)]$ can be computed in an analogous way. ∎
In the next lemma, we will demonstrate that it suffices to assume $x_{1}=x_{2}$, whenever we want to minimize $T_{\max}(G)$.
Lemma 2.
If $G\in{\mathbb{F}}_{q}^{2\times n}$ is a matrix with $x_{1}>x_{2}$, then there exists a matrix $G^{\prime}\in{\mathbb{F}}_{q}^{2\times 2n}$ with $T_{\max}(G^{\prime})<T_{\max}(G)$.
Proof.
Assume that $x_{1}>x_{2}$. From Equation (3), we have
$T_{\max}(G)={\mathbb{E}}[\tau_{2}(G)]=1+\frac{x_{1}}{x-x_{1}}+\sum_{i=0}^{q-2}\frac{a_{i}}{x-a_{i}}.$
Now, define the concatenated matrix $G^{\prime}=G\circ\overline{G}\in{\mathbb{F}}_{q}^{2\times 2n}$, where $\overline{G}$ is a modified version of $G$, in which each column of the form $(1,0)^{T}$ is replaced by $(0,1)^{T}$, and vice versa.
For $G^{\prime}$, we have
$T_{\max}(G^{\prime})$ $=1+\frac{x_{1}+x_{2}}{2x-x_{1}-x_{2}}+\sum_{i=0}^{q-2}\frac{2a_{i}}{2x-2a_{i}}$
$=1+\frac{x_{1}+x_{2}}{2x-x_{1}-x_{2}}+\sum_{i=0}^{q-2}\frac{a_{i}}{x-a_{i}}.$
Since $x_{1}>x_{2}$, it follows that $T_{\max}(G^{\prime})<T_{\max}(G)$. This completes the proof of Lemma 2. ∎
The next lemma is analogous to Lemma 2 but for the $a_{i}$ values, i.e., the number of columns of the form $(1,\beta^{i})^{T}$ in $G$.
Lemma 3.
If $G\in{\mathbb{F}}_{q}^{2\times n}$ is a matrix with $a_{i}<a_{j}$ for some $i,j\in\{0,1,\dots,q-2\}$, then there exits a matrix $G^{\prime}\in{\mathbb{F}}_{q}^{2\times 2n}$ with $T_{\max}(G^{\prime})<T_{\max}(G)$.
Proof.
Similarly to the proof of Lemma 2, define a matrix $G^{\prime}=G\circ\overline{G}\in{\mathbb{F}}_{q}^{2\times 2n}$, where $\overline{G}$ is a modified version of $G$, in which each column of the form $(1,\alpha^{i})^{T}$ is replaced by $(1,\alpha^{j})^{T}$, and vice versa. Then we have
$T_{\max}(G^{\prime})-T_{\max}(G)=\frac{2a_{i}+2a_{j}}{2x-a_{i}-a_{j}}-\frac{a_{i}}{x-a_{i}}-\frac{a_{j}}{x-a_{j}}=x(-a_{i}^{2}-a_{j}^{2}+2a_{i}a_{j}).$
Thus, since $a_{i}\neq a_{j}$, we have that $T_{\max}(G^{\prime})-T_{\max}(G)<0$. ∎
By Lemmas 2 and 3 we can assume that in an optimal matrix $G$, $x_{1}=x_{2}$ and $a_{i}=a_{j}$ for all $i,j\in\{0,\dots,q-2\}$. Substituting these assumptions in Equation (3) (let $a:=a_{i}$ and so $x=2x_{1}+(q-1)a$), we obtain
$T_{\max}(G)={\mathbb{E}}[\tau_{1}(G)]=1+\frac{x_{1}}{x_{1}+(q-1)a}+\frac{(q-1)a}{2x_{1}+(q-2)a}.$ (5)
In the next theorem, we determine the value of $T_{q}(2)$.
Theorem 1.
We have that $T_{q}(2)=1+\frac{2q^{2}-q(\sqrt{2}+1)-2+\sqrt{2}}{q^{2}(1+\sqrt{2})-q(2+\sqrt{2})}$.
Proof.
We begin by finding the optimal value of $a$ as a function of $x_{1}$. Taking the derivative of Equation (5) with respect to $a$ and setting the derivative to zero yields the condition
$-\frac{x_{1}(q-1)}{(x_{1}+(q-1)a)^{2}}+\frac{(q-1)(2x_{1}+(q-2)a)-(q-1)(q-2)a}{(2x_{1}+(q-2)a)^{2}}=0.$ (6)
Solving Equation (6),we find that the optimal value of $a$ is given by
$a=\frac{\sqrt{2}q-2}{q^{2}-2}x_{1}.$ (7)
This value corresponds to a minimum for Equation (5), since the derivative is negative before this value and positive afterwards. Note that although the value of $a$ is not an integer, due to its continuous dependence on $x_{1}$, it is always possible to select $x_{1}$ such that the value of $a$ in Equation (7) is arbitrarily close to an integer.
Substituting the value of $a$ from Equation (7) into Equation (5), we find
$T_{q}(2)$ $=1+\frac{q^{2}-2}{q^{2}-2+(q-1)(\sqrt{2}q-2)}+\frac{q-1}{\frac{2(q^{2}-2)}{\sqrt{2}q-2}+q-2}$
$=1+\frac{q^{2}-2}{q^{2}(1+\sqrt{2})-q(2+\sqrt{2})}+\frac{\sqrt{2}q^{2}-q(2+\sqrt{2})+2}{q^{2}(2+\sqrt{2})-q(2+2\sqrt{2})}$
$=1+\frac{q^{2}-2}{q^{2}(1+\sqrt{2})-q(2+\sqrt{2})}+\frac{q^{2}-q(\sqrt{2}+1)+\sqrt{2}}{q^{2}(\sqrt{2}+1)-q(\sqrt{2}+2)}$
$=1+\frac{2q^{2}-q(\sqrt{2}+1)-2+\sqrt{2}}{q^{2}(1+\sqrt{2})-q(2+\sqrt{2})}.\qed$
By taking $q$ to infinity we obtain the following corollary, which proves that the construction presented in [4, Theorem 12] is optimal.
Corollary 1.
It holds that $T(2)=\liminf_{q\longrightarrow\infty}T_{q}(2)=1+\frac{2}{\sqrt{2}+1}\approx 0.914\cdot 2$.
4 General Construction
In this section, we construct a $k$-dimensional code $\mathcal{C}_{k}$, which generalizes the construction presented in [17] specific to $k=3$. The construction will be presented in terms of a generator matrix for $\mathcal{C}_{k}$. In our construction, all columns of the generator matrix of the code, which we denote by $G_{k}$, are of weight $1$ or $2$. Consequently, the columns of $G_{k}$ can be partitioned into $\binom{k}{2}+k$ disjoint sets, where each set of columns corresponds to either a vertex or an edge in $K_{k}$, the complete graph on $k$ vertices. More precisely, we label $K_{k}=(V_{k},E_{k})$ in the following way:
  • (i)
    the set of vertices $V_{k}$ of $K_{k}$ correspond to the $k$ basis vectors $\textbf{e}_{i}$, $i\in[k]$, and we label by $V_{i}$ the vertex corresponding to $\textbf{e}_{i}$;
  • (ii)
    an edge between $\textbf{e}_{i_{1}}$ and $\textbf{e}_{i_{2}}$ for $i_{1},i_{2}\in[k]$ corresponds to the set of columns of $G_{k}$ with support $\{i_{1},i_{2}\}$ and we denote by $E_{i,j}$ the set of columns with support $\{i,j\}$.
Our objective is to ensure that $G_{k}$ is recovery complete, which we formally define as follows:
Definition 1.
A matrix $G_{k}$ is called recovery complete if the following conditions hold:
  1. 1.
    For any $j_{1},j_{2}\in[k]$, given two distinct columns ${{\boldsymbol{g}}}_{i},{{\boldsymbol{g}}}_{j}\in E_{j_{1},j_{2}}$, it is possible to recover the information strands $j_{1}$ and $j_{2}$, i.e., $\{{{\boldsymbol{e}}}_{i},{{\boldsymbol{e}}}_{j}\}\subseteq\langle{{\boldsymbol{g}}}_{i},{{\boldsymbol{g}}}_{j}\rangle$.
  2. 2.
    Suppose we collect one column each from $E_{j_{1},j_{2}},E_{j_{2},j_{3}},\ldots E_{j_{m},j_{1}}$ where the edges $\{j_{1},j_{2}\},\ldots,\{j_{m},j_{1}\}$ form a cycle in $K_{k}$ of length $m\leq k$. Then, it is possible to recover the information strands corresponding to the indices ${j_{1}},\ldots,{j_{m}}$.
We now turn to constructing recovery complete matrices, and we will begin by revisiting the case where $k=3$. While a construction for this case, based on a geometric viewpoint, was presented in [17], the approach taken here is designed so that it can be applied to a more general framework. Our generator matrices for the case where $k=3$ will have the following form:
[…]
where $G_{3}(x)\in{\mathbb{F}}_{q}^{3\times(3+3x)}$ and $\beta$ is a primitive element of ${\mathbb{F}}_{q}$.
Our objective is to show that for large enough $q$, we can find $\{i_{1},\dots,i_{3x}\}\subseteq\{0,\dots,q-2\}$ such that $G_{3}(x)$ is recovery complete, with each $E_{i,j}$ containing $x$ columns. To simplify, we assume $\mathbb{F}_{q}$ is a field of characteristic 2, allowing us to disregard signs.
To verify the first recovery completeness property, we check that any two columns in $E_{i,j}$ are linearly independent. For any two columns retrieved from $E_{i,j}$, the corresponding submatrix formed by their non-zero rows has the form:
$\begin{bmatrix}1&1\\ \beta^{i_{j}}&\beta^{i_{j^{\prime}}}\end{bmatrix}.$
Since the determinant of this matrix is non-zero whenever $i_{j}\not\equiv i_{j^{\prime}}\bmod(q-1)$, the result follows.
For the second property of recovery completeness, consider a matrix formed by selecting one column each from $E_{1,2},E_{1,3}$, and $E_{2,3}$. Note that such a matrix has the following form (up to permutation on the indices):
$\begin{bmatrix}1&0&1\\ \beta^{r}&1&0\\ 0&\beta^{\ell}&\beta^{s}\end{bmatrix}.$
The determinant of this matrix is $\beta^{s}+\beta^{r+\ell}$, therefore, to ensure this determinant is non-zero, it suffices to ensure that $\beta^{s}\neq\beta^{r+\ell}$ or equivalently (as $\mathbb{F}_{q}$ is a field of characteristic 2) that
$s\not\equiv r+\ell\bmod(q-1).$ (12)
The requirement in (12) is identical to the construction of sum-free sets [11, 15]; A sum-free set $\mathcal{S}$ is defined as a set for which there are no solutions to the equation $r+\ell=s$ with $r,\ell,s\in\mathcal{S}$. For the setup where the sum-free set is defined over an abelian group of order $n$ (in our case $n=q-1$), it is known that if $n\not\equiv 0\bmod 2$ (as is the case here, since $q$ is assumed to be a power of $2$), then there exists a sum-free set of size at least $n(1/3-\frac{1}{3n})$. Let $\mathcal{S}\subseteq\{0,\dots,q-2\}$ be a sum-free set of size $x$. Setting the powers of $\beta$ in $E_{1,2}$, $E_{1,3}$, and $E_{2,3}$ to be equal to the elements of $\mathcal{S}$, yields the desired result. Although the requirement $x\leq\frac{q-1}{2}$ in [17], which is optimal, is superior to ours, our construction still ensures $q=\mathcal{O}(x)$, demonstrating feasibility.
In Construction 1, we extend these ideas to the case where $k>3$.
Construction 1.
Let $\mathcal{S}$ be a set of size at least $x\binom{k}{2}$ with the following property: for any $k^{\prime}\leq k$ distinct elements $i_{1},i_{2},\ldots,i_{k^{\prime}}\in\mathcal{S}$ (i.e., $i_{j}\neq i_{j^{\prime}}$ for $j\neq j^{\prime}$), the following holds:
$\ell_{1}i_{1}+\ell_{2}i_{2}+\cdots+\ell_{k^{\prime}}i_{k^{\prime}}\not\equiv 0\bmod(q-1),$ (13)
where $\ell_{i}\in\{-1,1\}$ for all $i\in[k^{\prime}]$, with at least one $\ell_{i}$ positive and another negative.
The generator matrix for our code is then given by:
$G_{k}(x)=\left[\begin{array}[]{cccccccccc}I_{k}&E_{1,2}&E_{1,3}&\cdots&E_{1,k}&E_{2,3}&\cdots&E_{2,k}&\cdots&E_{k-1,k}\end{array}\right]$ (15)
where there are $\binom{k}{2}$ sub-matrices $E_{i,j}$, each corresponding to a pair of distinct elements from $[k]$. Each submatrix $E_{i,j}$ has the following properties:
  1. 1.
    $|E_{i,j}|=x,$
  2. 2.
    the support of each column in $E_{i,j}$ is $\{i,j\}$,
  3. 3.
    the leading coefficient of each column in $E_{i,j}$ is $1$,
  4. 4.
    the second coefficient in each column of $E_{i,j}$ is a power of a fixed primitive element $\beta$, where the powers of $\beta$ in $E_{1,2}$ are the first $x$ elements of $\mathcal{S}$, the powers of $\beta$ in $E_{1,3}$ are the next $x$ elements of $\mathcal{S}$, and so on.
Construction 1 relies on the existence of so-called $B_{k-1}$ sequences over $\mathbb{Z}_{q-1}$, which are sequences satisfying the properties given in the following theorem. Note that we restate the theorem in the language of this paper.
Theorem 2 (see [5, Theorem 2]).
Let $q=2^{tk}$ for some $t\geq 1$. Then there exist $12^{t}+$ 1 integers $j_{0}=0,j_{1}=1,j_{2},\dots,j_{2^{t}+1}\leq q-2$ such that the sums
$j_{i_{1}}+\dots+j_{i_{k-1}}\quad(0\leq i_{1}\leq\dots\leq i_{k-1}\leq 2^{t})$
are all different mod $q-1$.
Lemma 4.
For $q=2^{tk}$, where $t=\lceil\log_{2}(x\binom{k}{2}-1)\rceil$, there exists a set $\mathcal{S}$ as described in Construction 1, and the generator matrix $G_{k}(x)$ in (15) is recovery complete.
Proof.
We begin by constructing a set $\mathcal{S}$ as described in Construction 1. Let $t$ be the smallest integer such that $x\binom{k}{2}\leq 2^{t}+1$. From Theorem 2 there exists a $B_{k-1}$ sequence over $\mathbb{Z}_{q-1}$ of size $12^{t}+$ 1 whenever $q=2^{tk}$, and we can take a subset $\mathcal{S}$ of this sequence of size $x\binom{k}{2}$ with $0\in\mathcal{S}$. Now if we have $k^{\prime}$ different elements $\{i_{1},\dots,i_{k^{\prime}}\}\in\mathcal{S}$ for which Equation (13) holds with equality, we can split the sum among the terms corresponding to positive $\ell_{i}$’s, and negative $\ell_{i}$’s, respectively, resulting in two different sums of (at most) $k-1$ terms which are equal mod $q-1$. Then we can add the appropriate number of 0’s to both sums (since $0\in\mathcal{S}$), so that both sums have exactly $k-1$ terms, and we get a contradiction since $\mathcal{S}$ is a $B_{k-1}$ sequence over $\mathbb{Z}_{q-1}$.
Next, we show that $G_{k}(x)$ is recovery complete. The fact that two columns from $E_{i,j}$ have full rank follows from the same reasoning as in the case of $k=3$.
For the second condition of recovery completeness, consider the scenario where we select one column each from $E_{i_{1},i_{2}},E_{i_{2},i_{3}},$ $\dots,E_{i_{m},i_{1}},$ and where $\{i_{1},i_{2}\},\{i_{2},i_{3}\}$, $\ldots$, $\{i_{m},i_{1}\}$ is a cycle of length $m\leq k$. To satisfy the second property, we must show that the following matrix has full rank:
$\begin{bmatrix}1&0&0&\cdots&0&1\\ \beta^{i_{1}}&a_{i_{2}}&0&\cdots&0&0\\ 0&b_{i_{2}}&a_{i_{3}}&\cdots&0&0\\ 0&0&b_{i_{3}}&\cdots&0&0\\ &&&\ddots&&\\ 0&0&0&\cdots&a_{i_{m-1}}&0\\ 0&0&0&\cdots&b_{i_{m-1}}&\beta^{i_{m}}\end{bmatrix},$ (16)
where in every column $1<j<m$, one of the elements is $1$ and the other is a power of $\beta$ of the form $\beta^{i_{j}}$, $i_{j}\in\mathcal{S}$. Since our underlying field has characteristic $2$, the determinant of this matrix is given by:
$\beta^{i_{m}}\prod_{j=2}^{m-1}a_{i_{j}}+\beta^{i_{1}}\prod_{j=2}^{m-1}b_{i_{j}}.$
Let $T\subseteq\{2,3,\dots,m-1\}$ denote the subset such that $a_{i_{j}}=1$ for every $j\in T$ and $a_{i_{j}}=\beta^{i_{j}}$ for every $j\notin T$. Then the determinant can be expressed as:
$\beta^{i_{m}+\sum_{j\notin T}i_{j}}+\beta^{i_{1}+\sum_{j\in T}i_{j}}.$
To ensure the determinant is non-zero, it suffices to require that:
$i_{m}+\sum_{j\notin T}i_{j}\not\equiv i_{1}+\sum_{j\in T}i_{j}\bmod q-1,$
where $i_{j}\neq i_{j^{\prime}}$ for $j\neq j^{\prime}$. This condition follows directly from property (13), ensuring that the matrix in (16) has full rank. Therefore, the generator matrix $G_{k}(x)$ from (15) is recovery complete. ∎
Observe that repeating $I_{k}$ multiple times in $G_{k}(x)$ does not affect the matrix’s recovery completeness. Henceforth, let $G_{k}(x,y)$ denote the matrix obtained from $G_{k}(x)$ by repeating $I_{k}$ $y$ times.
Before continuing with the analysis, we take a closer look at the structure of the recovery sets for recovery complete matrices. Similar to the discussion in Definition 1, we will find it useful to interpret the recovery process as one that is performed over the complete graph $K_{k}$ rather than over the columns of the matrix $G_{k}(x,y)$. More precisely, for every column, say $\bf g$, that is collected from $G_{k}(x,y)$, one of the following analogous events occurs over $K_{k}$:
  1. 1.
    If the column $\bf g$ has weight $1$ in row $j$, then we have collected the vertex labeled $V_{j}$ in the graph $K_{k}$.
  2. 2.
    If the column $\bf g$ has weight $2$ with support $\{j_{1},j_{2}\}$, then we have collected the edge $E_{j_{1},j_{2}}$ in $K_{k}$.
For shorthand, we will refer to the multiset of collected edges and vertices from $K_{k}$ using the notation $\mathscr{T}$.
Example 2.
For the case of $k=3$, the top left graph in Figure 1 represents the setup where the column with a single 1 in the first position is collected so that $\mathscr{T}=\{V_{1}\}$. The graph in the bottom left represents the event where three columns are collected where one column has support $\{1,2\}$, another column has support $\{2,3\}$, and a third column has support $\{1,3\}$ so that $\mathscr{T}=\{E_{1,2},E_{2,3},E_{1,3}\}$.
Figure 1: Recovery sets over $K_{3}$ for information strand $1$.
For convenience, we will refer to the scenario where a vertex is collected (i.e., where $V_{j}\in\mathscr{T}$) as a cycle of length $1$, and when an edge is collected twice as a cycle of length $2$. Using this interpretation, we can more precisely state conditions that allow for the recovery of any information strand in $G_{k}(x,y)$ or equivalently any vertex in $K_{k}$.
Claim 2.
The information strand corresponding to the vertex $V_{i}\in K_{k}$ can be recovered if and only if $V_{i}$ belongs to a connected component in $\mathscr{T}$ that contains a cycle.
Note that in the previous claim, we say that $V_{i}$ and $V_{j}$ belong to the same connected component if there exists a path from $V_{i}$ to $V_{j}$ that traverses only edges contained in the multiset $\mathscr{T}$. It is straightforward to verify that for each of the $6$ scenarios depicted in Figure 1, $V_{1}$ belongs to a connected component with at least one cycle and similarly for Figure 2. Note that the scenarios in Figure 2 are not all of the possibilities for recovering $V_{1}$ in $K_{4}$.
Figure 2: Recovery sets over $K_{4}$ for information strand $1$.
For general $k$, $x$, and $y$, it seems to be hard to explicitly compute $T_{\max}(G_{k}({x,y}))$. However, in [17] the authors computed $\lim_{x\to\infty}T_{\max}(G_{3}({x,x}))$.
4.1 Three Information Strands
In [17], the authors provided a construction of a generator matrix that is recovery complete for $k=3$. They provided an analysis of the asymptotic behavior of the random access expectation of a recovery complete generator matrix, for the case where the multiplicity of the information strands was the same as the one of strands per edge (i.e., $y=x$). They showed that the random access expectation in this case is bounded from above by $0.8822\cdot 3$. However, if we allow the multiplicity to be any value proportional to $x$ (i.e., $y=\alpha x$) we can achieve a strictly smaller value for the expectation, as we show below. The analysis follows the same steps as in [17], and the details can be found in Appendix A.
Theorem 3.
It holds that $\lim_{x\longrightarrow{}\infty}T_{\max}(G_{3}({x,\alpha x}))\leq 3-\frac{\alpha}{3+3\alpha}-\frac{2+10\alpha+5\alpha^{2}}{9(1+\alpha)^{2}}+\frac{(1+2\alpha)^{2}(1+2\alpha)}{9(1+\alpha)^{2}(2+\alpha)}+\frac{2\alpha^{2}(9+7\alpha)}{9(1+\alpha)^{2}(3+2\alpha)^{2}}.$
By numerical analysis we obtain that the value of $\alpha$ for which the minimum is attained is approximately $0.834$. By plugging $\alpha=0.834$ to Theorem 3 we obtain the following corollary, which slightly improves the result from [17].
Corollary 2.
$T_{\max}(G_{3}({x,0.834x}))\leq 2.645=0.881\overline{66}\cdot 3.$
4.2 Four Information Strands
In this subsection, we analyze the value of $T_{\max}(G_{4}({x,y}))$ using Lemma 1. In order to apply Lemma 1, we first need to find the values of $\alpha_{i}^{s}(G_{4}({x,y}))$ for all $s\in[6x+4y-1]$ and $i\in[k]$.
Lemma 5.
For all $i\in[k]$ we have
  • (i)
    $\alpha_{i}^{1}(G_{4}({x,y}))=y$;
  • (ii)
    $\alpha_{i}^{2}(G_{4}({x,y}))=3({x\choose 2}+x\cdot y)+{y\choose 2}+y(6x+3y)$;
  • (iii)
    $\alpha_{i}^{3}(G_{4}({x,y}))=3\left({3x+2y\choose 3}-{x+2y\choose 3}\right)-3({x\choose 3}+{x\choose 2}y+x{y\choose 2})+3x{x+2y\choose 2}+y{6x+3y\choose 2}+{y\choose 2}(6x+3y)+{y\choose 3};$
  • (iv)
    $\alpha_{i}^{s}(G_{4}({x,y}))={6x+4\choose s}-{3x+3y\choose s}-{x+2y\choose s-1}$  for $4\leq s\leq 3x+3y$;
  • (v)
    $\alpha_{i}^{s}(G_{4}({x,y}))={6x+4y\choose s}$  for $13x+3y+1\leq s\leq 6x+4y-$ 1.
Proof.
Suppose we want to recover ${\bf e}_{1}$.
(i) is easy to see.
In order to prove (ii), we need to count the number of 2-sets that contain ${\bf e}_{1}$ in their span. If none of the two columns is ${\bf e}_{1}$, then to recover ${\bf e}_{1}$, the 2-set must span an edge of the form $E_{1,j}$, and there are $3({x\choose 2}+x\cdot y)$ such 2-sets. Moreover, there are a total of (${y\choose 2}+y(6x+3y)$ $2$-sets that contain ${\bf e}_{1}$.
For (iii) we start by counting 3-sets that do not contain ${\bf e}_{1}$. Consider a $3$-set which lies in the span of $\langle{\bf e}_{1},{\bf e}_{j},{\bf e}_{j^{\prime}}\rangle$ for some $j,j^{\prime}\in\{2,3,4\}$. The only such $3$-sets that do not recover ${\bf e}_{1}$ are the ones where all the columns lie in $\langle{\bf e}_{j},{\bf e}_{j^{\prime}}\rangle$ and there are (${x+2y\choose 3}$ such 3-sets. Note that in this way, every such $3$-set which lies in $\langle{\bf e}_{1},{\bf e}_{j}\rangle$ was counted twice. For example if $j=2$, then the 3-sets contained in $\langle{\bf e}_{1},{\bf e}_{2}\rangle$ were counted as part of the 3-sets in $\langle{\bf e}_{1},{\bf e}_{2},{\bf e}_{3}\rangle$ but also as part of the 3-sets in $\langle{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\rangle$. There are (${x\choose 3}+{x\choose 2}y+x{y\choose 2}$ such sets (where we counted them depending on how many ${\bf e}_{2}$ they contain, which is 1, 2, or 3). The other $3$-sets that retrieve ${\bf e}_{1}$ but do not contain ${\bf e}_{1}$ are the sets where two columns lie in $\langle{\bf e}_{1},{\bf e}_{j}\rangle\setminus\{{\bf e}_{j}\}$ and the other column lies in $\langle{\bf e}_{i},{\bf e}_{i^{\prime}}\rangle$ for $i,i^{\prime}\neq j$, and there are $3x{x+2y\choose 2}$ such 3-sets. We are left to add the number of $3$-sets which contain ${\bf e}_{1}$ and there are $y{6x+3y\choose 2}+{y\choose 2}(6x+3y)+{y\choose 3}$ of them, where we counted them depending on how many ${\bf e}_{1}$ they contain, which is 1, 2, or 3.
(iv) We count the number of $s$-sets that do not recover ${\bf e}_{1}$. If the first coordinate in all the columns of the $s$-set is 00, then it does not recover ${\bf e}_{1}$, and there are (${3x+3y\choose s}$ such sets. The other case is when there is one column from $E_{1,i}$ for some $i\in\{2,3,4\}$, and the other $s-1$ columns must be contained in ${\bf e}_{j}\cup{\bf e}_{j^{\prime}}\cup E_{j,j^{\prime}}$ where $j,j^{\prime}\notin\{1,i\}$, and there are $3x{x+2y\choose s-1}$ such sets.
From the analysis of (iv), it is easy to verify that any $s$-set for $s\geq 3x+3y+1$ will recover ${\bf e}_{1}$.
All of the previous computations did not depend on ${\bf e}_{1}$, and thus we conclude that they are the same for all ${\bf e}_{i}$, $i\in[k]$. ∎
Despite having closed expressions for the $\alpha_{i}^{s}(G_{4}({x,y}))$ for all $s\in[6x+4y-1]$ and $i\in[k]$, and thus for $T_{\max}(G_{4}({x,y}))$ (see Lemma 1), it is challenging to compute the exact value of $T_{\max}(G_{4}({x,y}))$. Nonetheless, from experiments with the computer algebra program magma, we obtain the plot shown in Figure 3, as well as the result stated in Proposition 1.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.41.61.61.61.61.81.81.81.822220.850.850.850.850.870.870.870.870.890.890.890.890.910.910.910.910.930.930.930.93Ratio $\alpha=y/x$$x=5$$x=10$$x=100$$x=1000$
Figure 3: Normalized (by $k=4$) random access expectation $T_{\max}(G_{4}({x,y}))$ for various $x$ and ratios $\alpha=y/x$.
Proposition 1.
We have $T(4)\leq 0.851534\cdot 4$.
5 Conclusion and Future Research
In this paper, we addressed the random access problem in DNA storage. We determined the exact values of $T_{q}(2)$ and $T(2)$ and we extended the construction presented in [17]. Our analysis demonstrated that this generalization achieves the best-known results to date for $k=3$ and $k=4$. Despite these advancements, several intriguing questions remain open, which we aim to explore in future work. A couple of these directions are the following:
  1. 1.
    We have shown that the construction in Section 4 performs well for $k=3$ and $k=4$ and in particular the random access expectation of the construction for $k=4$ is smaller than $k=3$. Does this expectation continue to decrease as $k$ increases, and if so, what is its limit? Additionally, From [4], we know that for positive integers $a$ and $k$, it holds that $\frac{ak}{ak}\leq\frac{T(k)}{k}$, and we wonder whether it holds that $\frac{T(k+1)}{k+1}<\frac{T(k)}{k}$ for all $k$? If that is true, what is the limit of $\frac{T(k)}{k}$ as $k$ tends to infinity?
  2. 2.
    In all previous constructions of matrices achieving a low random access expectation, the number of columns of weight $m$ was the same for every set of $m$ coordinates. In other words, the marginal distribution of vectors of weight $m$ was uniform. However, it had not been formally proven that this uniformity is the optimal structure for minimizing $T_{\max}(G)$. For $k=2$, we have shown that the optimal matrix adheres to this uniform structure, but does this hold true for general $k$?
Appendix A Appendix
For the reader’s convenience, we restate the statement of the theorem before presenting the corresponding proof.
See 3
Proof.
By Corollary 4.8 in [17] we have:
$T_{\max}(G_{3}({x,\alpha x}))$ $=3+\frac{2}{3x(1+\alpha)-2}-\frac{\alpha x-1}{3x(1+\alpha)-1}-\frac{2(\alpha x^{2}+{x\choose 2})+\alpha x(3x+2\alpha x)+\frac{1}{2}\alpha x(\alpha x-1)}{{{3x(1+\alpha)-1}\choose{2}}}$
$+\sum_{s=3}^{x+2\alpha x}\prod_{i=0}^{s-1}\frac{x+2\alpha x-i}{3x(1+\alpha)-i-1}+\sum_{s=3}^{\alpha x+1}\frac{2x{\alpha x\choose{s-1}}}{{3x(1+\alpha)-1\choose s}}.$
We are going to evaluate each term separately. First:
$\lim_{x\xrightarrow{}\infty}\frac{2}{3x(1+\alpha)-2}=0.$
Second:
$\lim_{x\xrightarrow{}\infty}\frac{\alpha x-1}{3x(1+\alpha)-1}=\frac{\alpha}{3+3\alpha}.$
Third:
$\frac{2(\alpha x^{2}+{x\choose 2})+\alpha x(3x+2\alpha x)+\frac{1}{2}\alpha x(\alpha x-1)}{{{3x(1+\alpha)-1}\choose{2}}}$ $=\frac{2\alpha x^{2}+x^{2}-x+3\alpha x^{2}+2\alpha^{2}x^{2}+\frac{1}{2}\alpha^{2}x^{2}-\frac{1}{2}\alpha x}{\frac{1}{2}(3x(1+\alpha)-1)(3x(1+\alpha)-2)}$
$=\frac{(1+5\alpha+2.5\alpha^{2})x^{2}-(1+0.5\alpha)x}{\frac{1}{2}(3x(1+\alpha)-1)(3x(1+\alpha)-2)},$
which goes to $\frac{2+10\alpha+5\alpha^{2}}{9(1+\alpha)^{2}}$ when $x$ goes to infinity.
Next:
$\sum_{s=3}^{x+2\alpha x}\prod_{i=0}^{s-1}\frac{x+2\alpha x-i}{3x(1+\alpha)-i-1}=\sum_{s=3}^{x+2\alpha x}\frac{x+2\alpha x}{3x(1+\alpha)-1}\cdot\frac{x+2\alpha x-1}{3x(1+\alpha)-2}\prod_{i=2}^{s-1}\frac{x+2\alpha x-i}{3x(1+\alpha)-i-1}$
Since for $i\geq 2$ we have that $\frac{x+2\alpha x-i}{3x(1+\alpha)-i-1}\leq\frac{1+2\alpha}{3+3\alpha}$, we obtain that
$\sum_{s=3}^{x+2\alpha x}\prod_{i=0}^{s-1}\frac{x+2\alpha x-i}{3x(1+\alpha)-i-1}\leq\frac{x+2\alpha x}{3x(1+\alpha)-1}\cdot\frac{x+2\alpha x-1}{3x(1+\alpha)-2}\sum_{s=3}^{x+2\alpha x}(\frac{1+2\alpha}{3+3\alpha})^{s-2}.$
Now we are going to evaluate $\sum_{s=3}^{x+2\alpha x}(\frac{1+2\alpha}{3+3\alpha})^{s-2}$. It holds that
$\sum_{s=3}^{x+2\alpha x}(\frac{1+2\alpha}{3+3\alpha})^{s-2}$ $=\sum_{s=1}^{x+2\alpha x-2}(\frac{1+2\alpha}{3+3\alpha})^{s}=\sum_{s=0}^{x+2\alpha x-2}(\frac{1+2\alpha}{3+3\alpha})^{s}-1$
$=\frac{(\frac{1+2\alpha}{3+3\alpha})^{x+2\alpha x-1}-1}{\frac{1+2\alpha}{3+3\alpha}-1}-1=\frac{(\frac{1+2\alpha}{3+3\alpha})^{x+2\alpha x-1}-1}{\frac{-2-\alpha}{3+3\alpha}}-1.$
Combining all of the above, we obtain that
$\lim_{x\xrightarrow{}\infty}\sum_{s=3}^{x+2\alpha x}\prod_{i=0}^{s-1}\frac{x+2\alpha x-i}{3x(1+\alpha)-i-1}$
$\leq\lim_{x\xrightarrow{}\infty}\frac{x+2\alpha x}{3x(1+\alpha)-1}\cdot\frac{x+2\alpha x-1}{3x(1+\alpha)-2}(\frac{(\frac{1+2\alpha}{3+3\alpha})^{x+2\alpha x-1}-1}{\frac{-2-\alpha}{3+3\alpha}}-1)$
$=\frac{1+2\alpha}{3(1+\alpha)}\cdot\frac{1+2\alpha}{3(1+\alpha)}\cdot(\frac{3+3\alpha}{2+\alpha}-1)=\frac{(1+2\alpha)^{2}}{9(1+\alpha)^{2}}\cdot(\frac{1+2\alpha}{2+\alpha}).$
And we are left with the last term.
$\sum_{s=3}^{\alpha x+1}\frac{2x{\alpha x\choose{s-1}}}{{3x(1+\alpha)-1\choose s}}$ $=^{(a)}2x\sum_{s=3}^{\alpha x+1}\frac{\frac{s}{\alpha x+1}{\alpha x+1\choose s}}{{3x(1+\alpha)-1\choose s}}=^{(b)}2x\sum_{s=3}^{\alpha x+1}\frac{s}{\alpha x+1}\frac{{3x(1+\alpha)-1-s\choose\alpha x+1-s}}{{3x(1+\alpha)-1\choose\alpha x+1}}$
$=\frac{2x}{\alpha x+1}{3x(1+\alpha)-1\choose\alpha x+1}^{-1}\sum_{s=3}^{\alpha x+1}s{3x(1+\alpha)-1-s\choose\alpha x+1-s},$
Where $(a)$ holds because (${\alpha x\choose{s-1}}=\frac{s}{\alpha x+1}{\alpha x+1\choose s}$ and $(b)$ is due to
(${\alpha x+1\choose s}{3x(1+\alpha)-1\choose\alpha x+1}={3x(1+\alpha)-1-s\choose\alpha x+1-s}{3x(1+\alpha)-1\choose s}.$
Now we are going to evaluate the term: $\sum_{s=3}^{\alpha x+1}s{3x(1+\alpha)-1-s\choose\alpha x+1-s}$, by placing $T=\alpha x+1-s$ we have that:
$\sum_{s=3}^{\alpha x+1}s{3x(1+\alpha)-1-s\choose\alpha x+1-s}$ $=\sum_{T=0}^{\alpha x-2}(\alpha x+1-T){3x(1+\alpha)-1+T-\alpha x-1\choose T}$
$=\sum_{T=0}^{\alpha x-2}(\alpha x+1){3x+2\alpha x-2+T\choose T}-\sum_{T=0}^{\alpha x-2}T{3x+2\alpha x-2+T\choose T}.$
Now, by applying Pascal’s triangle recursively on (${3x+3\alpha x-3\choose\alpha x-2}$ we have that
$\sum_{T=0}^{\alpha x-2}{3x+2\alpha x-2+T\choose T}={3x+3\alpha x-3\choose\alpha x-2}$
and hence we obtain:
$\sum_{T=0}^{\alpha x-2}(\alpha x+1){3x+2\alpha x-2+T\choose T}$ $=(\alpha x+1){3x+3\alpha x-3\choose\alpha x-2}$
$=(\alpha x+1){3x+3\alpha x-3\choose\alpha x-1}\frac{\alpha x-1}{3x+2\alpha x-1}.$
In addition, we have:
$T{3x+2\alpha x-2+T\choose T}$ $=T\frac{(3x+2\alpha x-2+T)!}{(T!)(3x+2\alpha x-2)!}$
$=(3x+2\alpha x-1)\frac{(3x+2\alpha x-2+T)!}{(T-1)!(3x+2\alpha x-1)!}$
$=(3x+2\alpha x-1){3x+2\alpha x-2+T\choose T-1}.$
Hence we can rewrite:
$\sum_{T=0}^{\alpha x-2}T{3x+2\alpha x-2+T\choose T}$ $=(3x+2\alpha x-1)\sum_{T=1}^{\alpha x-2}{3x+2\alpha x-2+T\choose T-1}$
$=(3x+2\alpha x-1)\sum_{T=0}^{\alpha x-3}{3x+2\alpha x-1+T\choose T}$
$=(3x+2\alpha x-1){3x+3\alpha x-3\choose\alpha x-3}$
$=\frac{(\alpha x-2)(\alpha x-1)}{3x+2\alpha x}{3x+3\alpha x-3\choose\alpha x-1}.$
Combining all of the above we obtain
$\sum_{s=3}^{\alpha x+1}\frac{2x{\alpha x\choose{s-1}}}{{3x(1+\alpha)-1\choose s}}$ $=\frac{2x}{\alpha x+1}{3x(1+\alpha)-1\choose\alpha x+1}^{-1}\Bigg{[}\frac{(\alpha x+1)(\alpha x-1)}{3x+2\alpha x-1}{3(x+\alpha x-1)\choose\alpha x-1}$
$\quad-\frac{(\alpha x-2)(\alpha x-1)}{3x+2\alpha x}{{3(x+\alpha x-1)\choose\alpha x-1}}\Bigg{]}$
$=\frac{2x(\alpha x-1)}{\alpha x+1}{3x(1+\alpha)-1\choose\alpha x+1}^{-1}{3(x+\alpha x-1)\choose\alpha x-1}\left[\frac{\alpha x+1}{3x+2\alpha x-1}-\frac{\alpha x-2}{3x+2\alpha x}\right]$
$=\frac{2x(\alpha x-1)}{\alpha x+1}\frac{(\alpha x+1)\alpha x}{(3x+3\alpha x-1)(3x+3\alpha x-2)}\left[\frac{9x+7\alpha x-2}{(3x+2\alpha x-1)(3x+2\alpha x)}\right]$
which goes to $\frac{2\alpha^{2}(9+7\alpha)}{9(1+\alpha^{2})(3+2\alpha)^{2}}$ when $x$ goes to infinity. ∎